OpenStudy (anonymous):

2x+5+4x=7x+5-x solve for equation if its an identity or no solution ?

OpenStudy (anonymous):

I came up with 2(0)+5+4(0)=7(0)+5-0 5=5 this would make the equation correct?

OpenStudy (anonymous):

THis would make it an identity.

OpenStudy (anonymous):

Identities are when it's like so: 6 = 6, 3 = 3, 0 = 0. No solutions are when it's like 0 = 4, 3 = 5, 4 = 8.